MONT 104Q – Mathematical Journeys Second Writing Assignment

The second of the larger paper assignments this semester will be devoted to the Adventures of Huckleberry Finn. This one will also lead to an argumentative essay in which you make a a particular response to a question – a thesis – and then present appropriate evidence and reasoning to support your claim(s). Your essay should show that you have done some careful reading, and some serious thinking, about the subject. You may, of course, disagree with ideas presented in class if you want. The point again is for you to develop your own thinking, and then clearly express it in writing.

Your essay should be 3-5 pages in length (double-spaced, single-sided). Please submit as a MS Word .doc or .docx file to jlittle@holycross.edu by 5:00pm on Monday, October 26.

For this assignment you have a choice of which one of the two following topics you will address:

1. Twain himself said that “Huck Finn is a book of mine where a sound heart and a deformed conscience come into collision and conscience suffers defeat.” What do you think he meant by a sound heart and a deformed conscience? There are several key stages in the story and important decisions made by Huck that show him going against the feelings of conscience he has acquired from his upbringing and his surroundings. What are those stages and decisions? When and how do they occur, and why specifically can we say that Huck’s conscience was somehow “deformed?” Does Huck change over the course of the book, and if so, how? And how does all of this connect with our CHQ cluster theme?

2. We know that Twain often recounted the final chapters of Huck Finn, beginning with Huck’s arrival at the Phelps farm, out loud from memory as a sort of humorous “set piece” in the speaking tour he undertook to publicize the book. Why do you suppose he chose these chapters to perform in that way? He says his performances were always received well by his audiences. But do these chapters seem out of place with the rest of the book to you? If so, why? What is it about them that seems different? Even if you don’t think they are out of place, if we see the end of Chapter XXXI as kind of a climactic moment, do these final chapters build on that climax or lessen its impact?

Guidelines for this assignment

Follow these specifications:

1. It is OK to discuss preliminary ideas for what you want to say with your classmates if you want. I will also be happy to talk over things and/or read a preliminary draft to give feedback if you like.

2. In the paper itself, develop, and then clearly state your thesis in your opening paragraph. As in the first paper this should be a clearly-articulated answer to one of the two topic questions posed above. The thesis should be clearly stated and well focused. The reader (me) should not need to guess what your point is.

3. In the following paragraphs, using sound reasoning, present your evidence for making the claim in your thesis. Do not contradict what you said earlier. Do not ignore evidence that contradicts your ideas. If, on further reflection, you find that you do not agree with what you said before, revise the thesis(!)

4. In this sort of writing, a “move” that is often very effective is to temporarily entertain an opposing conclusion as a possibility, then explain the evidence that leads you to reject that conclusion. Try it!

5. Your evidence should consist of passages from The Adventures of Huckleberry Finn, and interpretations of those passages that come from your own thinking about them. Your evidence should be as specific, detailed, correct, and as varied as possible (focus on small passages and mention the little things that support your point).

6. Do not consult any other sources beyond Twain’s novel – this means no other books, no web sites, etc.

7. When you use direct quotations, at the end place the Chapter and page number(s) to identify the source.

8. The paper as a whole should be well-organized and “flow” from one point to the next. Your sentences should be clear, readable, and not overly wordy. Aim for a reasonably formal but lively style. Careless or colloquial ways of saying things that are acceptable in casual conversation are not appropriate in this kind of writing. Also, please do not use a thesaurus to find impressive-sounding words when simpler alternatives are available. An important lesson to learn for writing in college is that thesauri do not capture the nuances in meanings that underlie good writing. If you must use a thesaurus, also look up any words you select in a dictionary to make sure they mean exactly what you think they do!

9. Errors in mechanics and usage should not annoy the reader. (As a warning, I tend to be easily annoyed by this sort of thing!) Try as much as possible to avoid mistakes in grammar, spelling, punctuation, and word choice.